OpenStudy (anonymous):

0.5(y+2)-0.3=0.4y+0.5

OpenStudy (anonymous):

0.5y +1 -0.3-0.4y-0.5=0 can you finish it?

OpenStudy (anonymous):

If I could I wouldnt have asked

OpenStudy (anonymous):

0.1y+0.2=0 y=-0.2/0.1 y=-2